[Brahms-l] Shift report 20040102 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Fri Jan 02 2004 - 23:59:51 EST
------------------------------------------------------------
Date: Fri Jan  2 23:59:50 2004, Shift: 16:00-24:00
Supervisor: sjs, with: 
Log book#15 pages 22 - 22

Shift Summary:
------------------------------------------------------------
1. No beam until 21:15.  Continuous store since then.
2. Pedestal (9326,9328-partial), TPC Pedestal(9327).  Initial 
pedestal run 9326 found to take over hour. I realized
afterwards that I by following our current instructions I was
missing the ZDCs. I subsequently started a second run (9328)
with FB1 selected, but stopped that short when we started to
get beam.
Data runs (trig 1):  9329(343K), 9331(77K), started 9332

3. No new detector problems found.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 3 00:00:16 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 03 2004 - 00:00:34 EST